# A Comparative Analysis of Large Language Models for Code Documentation Generation

### v1

@@ 4.3. Correlation between different metrics
Figure 8. Documentation level wise metrics comparison
Figure 9. Documentation level wise time taken comparison
@@ 4.4. Statistical Analysis
For statistical analysis, ANOVA (Analysis of Variance) (St»hle and Wold, 1989) is employed to assess the significance of differences among group means in multi-group comparisons. This method is particularly essential when evaluating multiple models across various parameters, such as performance metrics. ANOVA utilizes F-statistics and p-values to determine whether observed variances across models are statistically significant or occur by chance, thereby providing a rigorous and technical approach to model evaluation and comparison.
@@ 5. Conclusion
Following are the findings of our comparative analysis:
  • With the exception of Starchat, all Large Language Models (LLMs) exhibit either equal or superior performance when compared to the original documentation. Starchat consistently yielded suboptimal results when compared to the original documentation.
  • Closed-source models, including GPT-3.5, GPT-4, and Bard, consistently outperform their open-source counterparts, Llama2 and Starchat, across a majority of parameters in our evaluation rubric. Notably, when considering the time taken for generation, GPT-4 demonstrated the longest duration, followed by Llama2, Bard, with ChatGPT and Starchat having comparable generation times.
  • Additionally, file level documentation had a considerably worse performance across all parameters (except for time taken) as compared to inline and function level documentation.

### v2

@@ 4.3. Correlation between different metrics
Figure 8. Documentation level-wise metrics comparison
Figure 9. Documentation level-wise time taken comparison
4.4. Comparison across different documentation levels
After looking at and comparing the performances of each LLM across different metrics, we will now look at the general performance of all the LLMs across different documentation levels, namely inline level documentation, function level documentation and file level documentation.
@@ 4.5. Statistical Analysis
For statistical analysis, ANOVA (Analysis of Variance) (St»hle and Wold, 1989) is employed to assess the significance of differences among group means in multi-group comparisons. This method is particularly essential when evaluating multiple models across various parameters, such as performance metrics. ANOVA utilizes F-statistics and p-values to determine whether observed variances across models are statistically significant or occur by chance, thereby providing a rigorous and technical approach to model evaluation and comparison. The alpha level or significance level used for this analysis is 0.05.
@@
5. Discussion
Our comparative analysis of code documentation generated by various Large Language Models (LLMs) yielded several significant findings. Firstly, with the exception of Starchat, all LLMs demonstrated either equivalent or superior performance compared to the original documentation, highlighting their potential for automating documentation tasks. Secondly, closed-source models like GPT-3.5, GPT-4, and Bard consistently outperformed their open-source counterparts, Llama2 and Starchat, across a majority of evaluation parameters. Notably, GPT-4 exhibited the longest time taken for generation, followed by Llama2 and Bard, while ChatGPT and Starchat had comparable times. Lastly, file-level documentation consistently displayed poorer performance across all parameters (except for time taken) when compared to inline and function-level documentation.
Following are some implications, recommendations, and suggestions for the software development process in accordance with the findings of our paper:
  1. (1)
    Performance Discrepancies between LLMs: The varying performance of LLMs underscores the importance of carefully selecting the model based on specific project requirements and evaluation criteria. Software developers should consider conducting pilot tests with multiple LLMs to determine which model best suits their documentation needs.
  2. (2)
    Closed-Source vs. Open-Source Models: The superiority of closed-source models suggests potential benefits in investing in proprietary LLMs for documentation tasks.
  3. (3)
    Time Taken for generation considerations: The variation in generation time among LLMs highlights the importance of balancing performance with efficiency in documentation tasks. Developers should weigh the trade-offs between documentation quality and generation time when selecting LLMs for their projects. Research efforts should focus on developing strategies to optimize generation time without compromising documentation quality, enhancing the overall efficiency of LLM based documentation workflows.
  4. (4)
    File-Level Documentation Challenges: The comparatively inferior performance of file-level documentation emphasizes the need for alternative approaches to improve its effectiveness. Future research could explore innovative methods for enhancing the quality and utility of file-level documentation, such as incorporating contextual information or utilizing specialized LLM architectures fine-tuned for larger code structures.
@@ 6. Conclusion
Based on our comparative analysis, some very interesting findings have emerged. Except for Starchat, all Large Language Models (LLMs) demonstrate either equivalent or superior performance when compared to the original documentation, with Starchat consistently yielding suboptimal results. Secondly, closed-source models such as GPT-3.5, GPT-4, and Bard consistently outperform their open-source counterparts, Llama2 and Starchat, across a majority of evaluation parameters. Notably, in terms of generation time, GPT-4 exhibits the longest duration, followed by Llama2 and Bard, while ChatGPT and Starchat have comparable generation times. Lastly, file-level documentation displays significantly poorer performance across all parameters (except for time taken) as compared to inline and function-level documentation.